OpenStudy (anonymous):

Differentiate the following with respect to x by first principles: 3 cos (4x) − 5 sin (4x)

OpenStudy (anonymous):

\[f(x)\prime=\lim_{h \rightarrow 0} \frac{ f(x+h) -f(x)}{ h }\]

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

do you have to use this limit ?

OpenStudy (reemii):

what do you call "first principles' ?

OpenStudy (anonymous):

that's first principles formula (limit)

OpenStudy (anonymous):

we have to get rid of h at the denominator, then replace h with zero

OpenStudy (reemii):

you probably probably have somewhere in your course the proof, with this limit, that \((\sin(x))'=\cos( x)\). You should take this and use this as base for this question.

OpenStudy (amistre64):

you will need to remember your trig(a+b) identities

OpenStudy (amistre64):

3 cos (4(x+h)) − 5 sin (4(x+h)) 3 cos (4x+4h) − 5 sin (4x+4h) 3(cos(4x)cos(4h)-sin(4x)sin(4h)) - 5(sin(4x)cos(4h)+sin(4h)cos(4x)) .... etc

OpenStudy (amistre64):

and yes, recalling the sinh/h and cosh/h limits from squeeze thrm would have to be applied